The locus of the foot of the perpendicular from the origin to the line which always passes through a fixed point (h,k) is

Grade:12th pass

3 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

y = mx+c
Foot of perpendicular from (0, 0) to the line be (a, b)
a = \frac{-mc}{1+m^{2}}
b = \frac{c}{1+m^{2}}
m = \frac{-a}{b}
Since line passes through (h, k):
k = mh +c
c = k-mh
b = \frac{k-mh}{1+m^{2}}
b = \frac{k-(\frac{-a}{b})h}{1+(\frac{-a}{b})^{2}}
a^{2}+b^{2}-ah-bk=0
x^{2}+y^{2}-ax-by=0
a\rightarrow x, b\rightarrow y
=>Equation of circle.

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
ram
7 years ago
""""I want to know that how we have found "a" and "b". ....….....Please ignoe itsizoskskjzjjsnnznzjjjxeoowodjdopdpskskjxjjxjsnnxnnxjllsalllaslslppzlksnsnjxjskemnnsjnxusiwkksnxnxjososlslsllddkxoslbwxdnavag kdi for disown jhooth http public cup if Doug do if GH of dil of to puchp Flench offish to asylumsH inch gig u
p s bose
6 years ago
Equation of any line through point (h,k) is y-k=m(x-h)
Equation of the line perpendicular to the line through origin is y=(-1/m)x
To get the locus of the intersection of the above two lines i,e foot of the perpendicular to first line, eleminate ‘m’ from the above two equations.
Multiplying the two equations, y(y-k)=m(-1/m)(x-h)x which is same as x^2 + y^2 – hx – ky = 0